# Embodied Co-Design for Rapidly Evolving Agents: Taxonomy, Frontiers, and Challenges

### v1

@@ III-D Co-Design Constraints
Constraints are intrinsic to the design space of embodied agents. Unconstrained ECD often leads to overfitting: the morphological search may yield designs that are unnecessarily complex for a single task or, conversely, result in trivial solutions with poor transferability. Therefore, principled regularization is essential for guiding the search toward balanced brain-body pairs. Morphological constraints, such as structural stability, material availability, joint range-of-motion, and actuation limits, need to be considered during optimization. Common strategies include enforcing symmetry priors [dong2023symmetry, gupta2021embodied] and incorporating penalties for manufacturing time, cost, and complexity into objective functions, as demonstrated in monotone co-design [censi2015mathematical, zardini2021co, wilhelm2023constraint]. Behavioral constraints, such as safety considerations, actuator saturation, and collision avoidance, can further limit the feasible solution space and hinder the exploitation of simulators. Despite their significance, many ECD approaches still operate in largely unconstrained settings. Integrating practical engineering and fabrication constraints remains a challenge for the deployment of real-world applications.
@@ III-E Co-Design Frameworks
Based on the fundamental elements mentioned above, we further categorize ECD approaches into four frameworks: (1) bi-level co-design, (2) single-level co-design, (3) generative co-design, and (4) open-ended co-design. These categories reflect how algorithms optimize morphology, control, and, in some cases, the task environment. In Section IV, we elaborate on these frameworks and analyze the existing literature through the lens of our new taxonomy.
@@ III-F1 Task Performance
Task performance remains the primary application-level metric, directly measuring goal achievement. Evaluating this aspect is expensive during morphology search, because it typically requires complete policy optimization and extensive rollout-based testing.
@@ III-F2 Morphological Computation
which measures how informative the previous control state $s^{c}_{t}$ is about the next state $s^{c}_{t+1}$ when the action $a^{c}_{t}$ is known. A large MC indicates that morphology and environment strongly drive the state transition. While recent studies have directly quantified and optimized MC in physical robots [chandiramani2024improving, ghazi2016evaluating], practical applications often rely on simplified or surrogate estimators [zahedi2013quantifying, ghazi2015quantifying, fuchslin2013morphological, muller2017morphological].
@@ III-F3 Empowerment
Similar to MC, empowerment measures the maximum influence an agent can exert on future states through its actions [klyubin2005empowerment, de2018unified, leibfried2019unified, jung2011empowerment], and is widely used in reinforcement learning exploration [houthooft2016vime, kumar2018empowerment, choi2021variational]. Formally, the empowerment value for a control state $s^{c}_{t}$ is defined as the channel capacity between the action $a^{c}_{t}$ and the next state $s^{c}_{t+1}$:
@@
III-F4 Quality-Diversity (QD) Score
As discussed in Section IV-B4, diversity-based ECD approaches focus on discovering a wide repertoire of high-performing morphologies and behaviors. The QD score quantifies coverage and performance within a discretized Morphological Descriptor (MD) space, where each cell represents a morphology niche. For an archive with $N$ cells, the QD score is computed as the sum of the task returns $\mathcal{R}_{i}$ of the best individual in each occupied cell:
@@ III-F5 Complexity
Complexity measures the cost of both the controller and the morphology. Controller complexity can be quantified using architecture-agnostic metrics such as FLOPs, which estimate computational cost, latency, and energy consumption [xie2025morphology]. Morphological complexity can be quantified through component counts (e.g., links, joints, voxels, and degrees of freedom) or structural descriptors (e.g., connectivity, actuator distribution, material heterogeneity, and symmetry).
@@ III-F6 Robustness
Robustness evaluates an agent’s resilience to perturbations, including morphological damage (e.g., locked joints, sensor faults) and environmental variations. Robust agents maintain high performance and enable rapid adaptation, reflecting their inherent fault tolerance [powers2021good].
@@ IV Methods: Synthesizing Embodied Co-Design
Figure 4: An overview of four embodied co-design frameworks. Example images of co-designed agents from (a) to (e) are adapted from DERL [gupta2021embodied], Softzoo [wang2023softzoo], EvolutionGym [bhatia2021evolution], RoboGrammar [zhao2020robogrammar], and [ha2019reinforcement].
In this section, we elaborate on four ECD frameworks and organize them into sub-categories following the taxonomy outlined in Section III. For each sub-category, we analyze its core methodological components and highlight recent approaches. An overview of these frameworks is presented in Fig. 4.
@@ IV-A1 Evolutionary Reinforcement Learning
In this method, the inner loop for control policy optimization typically employs reinforcement learning (RL) [singh2022reinforcement, wang2022deep], whereas the outer loop for morphology evolution relies on various evolutionary algorithms (EAs).
@@
Two major classes of RL methods are used for policy optimization: model-based and model-free approaches. Model-free RL learns a policy directly from interactions, without explicitly modeling transition dynamics or reward functions. Representative algorithms include proximal policy optimization (PPO) [schulman2017proximal], trust region policy optimization (TRPO) [schulman2015trust], deep deterministic policy gradient (DDPG) [silver2014deterministic, lillicrap2015continuous], twin delayed DDPG [fujimoto2018addressing], and soft actor-critic (SAC) [haarnoja2018soft]. Model-based RL, in contrast, leverages a provided or learned dynamics model to refine the control policy. Representative frameworks include Mb-Mf [nagabandi2018neural], PETS [chua2018deep], GPS [levine2013guided], and PILCO [deisenroth2011pilco]. Model-based approaches are more sample-efficient but can struggle in complex domains where accurate dynamics modeling is harder than direct policy learning.
@@
EAs are effective for exploring high-dimensional morphology spaces. In the outer loop, evolutionary operators determine how new morphologies are generated, thereby shaping both the breadth and depth of exploration [carvalho2024role]. In discrete design spaces, the mutation operator typically adds or removes structural elements—e.g., joints, limbs, modules, kinematic branches—or modifies material properties. In continuous design spaces—including link lengths, joint stiffness, and mass distribution—Gaussian perturbations are commonly used. The crossover operator further expands the search space by recombining structural components or genotype encodings from multiple parents. Notably, generative models, such as Large Language Models (LLMs), have been used as high-level mutation operators to propose structured, human-interpretable edits based on prompts [song2025laser], further expanding the operator toolbox for morphology search.
@@
By alternating between individual-level controller learning and population-level morphology evolution, ERL methods mirror Darwinian evolution augmented with a Baldwinian learning mechanism [luo2022effects, li2023evaluation]. A representative example is Neural Graph Evolution (NGE) [wang2019neural], which evolves graph-structured, rigid-body morphologies while training universal GNN-based controllers via PPO. NGE demonstrates that GNN controllers can generalize across related body structures, thereby reducing the cost of controller retraining during morphology search. The scalability of ERL is further illustrated by Deep Evolutionary Reinforcement Learning (DERL) [gupta2021embodied]. In contrast to NGE, DERL employs individual controllers that are trained from scratch for each morphology. Coupled with increasingly challenging environments—ranging from flat-ground locomotion to uneven terrain and object manipulation—this framework enables the emergence of morphological intelligence, resulting in agents that exhibit increased robustness, adaptability, and behavioral diversity. Similar strategies appear in later works [park2021computational, bhatia2021evolution, chen2023evolving, wang2024learn, yue2025toward].
@@
As summarized in Table I, recent research has also expanded the classical ERL workflow along several dimensions. First, richer morphological representations have been introduced, including hyperbolic embeddings [dongleveraging2024] and text-to-mesh generation [ringel2025text2robot]. Second, advanced control-learning techniques, such as meta-RL [belmonte2022meta], transfer learning [harada2024lamarckian], imitation learning [hernandezfollowing], and policy-reuse [xie2025accelerating], have improved sample efficiency and generalization. A parallel line of work introduces novel morphological selection mechanisms, such as successive halving [katayama2023ga], multi-fidelity archiving [zhao2024morphological, nagiredla2024ecode], and multi-objective balancing [zhao2025multi], which can prevent premature convergence in morphology space [doncieux2014beyond, mertan2024investigating].
@@
Notably, the combination of EA and RL has also been widely explored in control optimization. These hybrid methods maintain a population of policies that evolve across generations, enabling broad exploration of the policy space while refining individual controllers via RL’s gradient-based or value-based updates. Representative approaches include ERL [khadka2018evolution], CERL [khadka2019collaborative], PDERL [bodnar2020proximal], SERL [wang2022surrogate], and ERL-Re2 [hao2023erl]. Comprehensive reviews of this line of work can be found in [drugan2019reinforcement, sigaud2023combining, bai2023evolutionary, zhu2023survey, lin2025evolutionary, li2024bridging].
IV-A2 Surrogate-Assisted Methods
A central challenge in bi-level co-design arises from its hierarchical structure, where the inner-loop optimization is computationally intensive and often requires days of training for complex tasks. Surrogate-assisted methods address this issue by employing predictive models to approximate task performance, thereby reducing the need for repeated full-scale policy evaluations.
Bayesian Optimization (BO) [shahriari2015taking] has emerged as a prominent framework for modeling the morphology–performance relationship in ECD. As a black-box, zeroth-order optimizer, BO efficiently handles expensive objective evaluations. At each iteration, BO updates a surrogate model using previously evaluated morphologies and their performance, then selects new candidates through acquisition functions such as Probability of Improvement (PI), Expected Improvement (EI), Entropy Search (ES), or Upper Confidence Bound (UCB), thereby balancing exploration and exploitation [wilson2018maximizing]. The selected candidates are evaluated, and their results are used to update the surrogate for the subsequent iteration.
Gaussian Processes (GPs) are among the most commonly used surrogates in BO. For instance, HPC-BBO [liao2019data] introduced a hierarchical framework that facilitates parallel hardware validation. Similarly, [bjelonic2023learning] pre-trained a universal controller in simulation and used BO to optimize parallel–elastic spring parameters for a real quadruped, demonstrating effective sim-to-real transfer. Variants of BO have been further explored in [bhatia2021evolution, hu2021neural, saito2024effective, pan2021emergent]. To enhance generalization, [chen2024pretraining] combined domain randomization with a pretraining–finetuning pipeline. Extending beyond single-task BO, [dong2024cagecoopt] proposed a multi-task BO framework for robot manipulators. More recently, [schneider2025task] combined BO with Hyperband to generate novel manipulator designs that improve performance across both seen and unseen tasks.
In addition to traditional surrogates used in BO, recent approaches have explored alternative choices. For example, Q-value functions (Q-Surro) from RL have been used to evaluate candidate morphologies [luck2020data, luck2021robot, du2025efficient]. Multi-dynamics models can further enhance generalization across different agent designs [sliacka2023co]. Alternatively, when historical interaction data is available, near-optimal control actions (A-Surro) can be partially or fully transferred from previously optimized designs to accelerate evaluation [liu2023rapidly]. More recent studies by [zhao2025cross, strgar2025accelerated] developed universal policies that span multiple tasks and can be efficiently used as a zero-shot performance evaluator (A-Surro), thereby improving efficiency. In [bai2025learning], manipulator morphologies are evaluated and optimized using pre-collected trajectories and a learned reward model (R-Model), with the reliance on reward model evaluations gradually increasing throughout the training process.
@@ IV-A2 Surrogate-Assisted Methods
Learnable (CL) Individual Discrete (UF) Direct Modifiable (Full) Task-Driven
@@ IV-A3 Other Methods
The inner-loop controller optimization can be implemented using alternatives to RL, including trajectory optimization [fadini2022simulation], evolutionary algorithms [jelisavcic2019lamarckian, le2022morpho, le2025efficient], numerical solvers [sartore2023codesign, ha2018computational2], and information-theoretic classifiers [iii2021taskagnostic]. To accelerate bi-level co-design, Goff et al. proposed A-MEL [goff2024efficient], which integrates asynchronous learning and evolution to reduce both time and computational costs. Strgar et al. [Strgar-RSS-24] further leveraged massively parallel differentiable simulations to efficiently optimize neural controllers for large populations of candidate morphologies.
@@ IV-B Single-Level Co-Design
As shown in Fig. 4 (b), single-level co-design treats both morphology and control as variables within a unified optimization process, unlike bi-level co-design, which employs nested loops for these aspects. We organize recent work along this line into four subcategories: 1) EA-Based Methods, 2) RL-Based Methods, 3) Physics-Based Methods, and 4) Diversity-Based Methods, which are summarized in Table III.

### v2

@@ III-D Co-Design Constraints
Despite their significance, many ECD approaches still operate in largely unconstrained settings, which can lead to overfitting. Specifically, morphological search may yield designs that are unnecessarily complex for a single task or, conversely, result in trivial solutions with poor transferability. Therefore, principled regularization is essential for guiding the search toward balanced brain-body pairs.
Morphological constraints, such as structural stability, material availability, joint range-of-motion, and actuation limits, can be explicitly incorporated into the design process. Common strategies include enforcing symmetry priors [dong2023symmetry, gupta2021embodied] and penalizing manufacturing time, cost, and complexity in the objective function, as in monotone co-design [censi2015mathematical, zardini2021co, wilhelm2023constraint]. Behavioral constraints, such as safety, actuator saturation, and collision avoidance, can further limit the feasible behavior space and hinder the overuse of simulators.
@@ III-E Co-Design Frameworks
Based on the fundamental elements mentioned above, we categorize ECD approaches into four frameworks—(1) bi-level co-design, (2) single-level co-design, (3) generative co-design, and (4) open-ended co-design—because these categories capture the most significant differences in how algorithms coordinate morphology, control, and, optionally, the task environment during optimization (Section IV).
The bi-level/single-level split distinguishes between hierarchical (nested) and unified (joint) optimization loops, thereby separating methods based on evaluation cost, convergence behavior, and morphological diversity. The generative category isolates approaches that rely on learned or rule-based priors to generate candidate bodies, emphasizing the trade-offs between sample efficiency and design freedom. Open-ended co-design, finally, groups methods that promote continual innovation through the co-evolution of task environments and agents. Together, these axes establish a design space for ECD methods that can guide the development of new embodied agents and their customization for specific applications.
@@ III-F1 Task Performance
Task performance remains the primary application-level metric, as it directly measures goal achievement. Evaluating this aspect is costly during morphology search because it typically requires complete policy optimization and extensive rollout-based testing.
@@ III-F2 Morphological Computation
which measures how informative the previous control state $s^{c}_{t}$ is about the next state $s^{c}_{t+1}$ when the action $a^{c}_{t}$ is known. A large MC indicates that morphology and environment strongly drive state transitions. Although recent studies have directly quantified and optimized MC in physical robots [chandiramani2024improving, ghazi2016evaluating], most practical applications rely on simplified or surrogate estimators [zahedi2013quantifying, muller2017morphological].
@@ III-F3 Empowerment
Similar to MC, empowerment measures the maximum influence an agent can exert on future states through its actions [klyubin2005empowerment, leibfried2019unified], and it is widely used in reinforcement learning exploration [houthooft2016vime]. Formally, the empowerment value for a control state $s^{c}_{t}$ is defined as the channel capacity between the action $a^{c}_{t}$ and the next state $s^{c}_{t+1}$:
@@
III-F4 Quality-Diversity Score
As discussed in Section IV-B4, diversity-based ECD methods aim to discover a wide repertoire of high-performing morphologies and behaviors. Accordingly, the Quality-Diversity (QD) score measures both coverage and performance over a discretized Morphological Descriptor (MD) space, where each cell corresponds to a distinct morphological niche. For an archive with $N$ cells, the QD score is computed as the sum of the task returns $\mathcal{R}_{i}$ of the best individual in each occupied cell:
@@ III-F5 Complexity
Complexity captures the cost associated with both the controller and the morphology. Controller complexity can be measured using architecture-agnostic metrics, such as FLOPs, which estimate computational load, latency, and energy consumption [xie2025morphology]. Morphological complexity can be quantified via component counts (e.g., links, joints, degrees of freedom) or structural descriptors (e.g., connectivity, material heterogeneity, and symmetry).
@@ III-F6 Robustness
Robustness evaluates an agent’s resilience to perturbations, including morphological damage (e.g., locked joints, sensor faults) and environmental variations [powers2021good]. Robust agents maintain high performance and enable rapid adaptation, reflecting their inherent fault tolerance.
@@ IV Methods: Synthesizing Embodied Co-Design
Fig. 3: An overview of four embodied co-design frameworks. Example images of co-designed agents from (a) to (e) are adapted from DERL [gupta2021embodied], Softzoo [wang2023softzoo], EvolutionGym [bhatia2021evolution], RoboGrammar [zhao2020robogrammar], and [ha2019reinforcement].
In this section, we elaborate on four ECD frameworks and organize them into sub-categories following the taxonomy outlined in Section III. For each sub-category, we analyze its core methodological components and highlight recent advances. An overview of these frameworks is presented in Fig. 3.
@@ IV-A1 Evolutionary Reinforcement Learning
In this method, the inner loop for control policy optimization typically employs reinforcement learning (RL) [singh2022reinforcement, wang2022deep], while the outer loop for morphology evolution utilizes various evolutionary algorithms (EAs).
@@
Two major classes of RL methods are used for policy optimization: model-based and model-free approaches. Model-free RL learns a policy directly from interactions without explicitly modeling transition dynamics. Representative algorithms include proximal policy optimization (PPO) [schulman2017proximal], trust region policy optimization (TRPO) [schulman2015trust], deep deterministic policy gradient (DDPG) [lillicrap2015continuous], twin delayed DDPG [fujimoto2018addressing], and soft actor-critic (SAC) [haarnoja2018soft]. In contrast, Model-based RL leverages either a provided or learned dynamics model to refine the control policy. Representative frameworks include Mb-Mf [nagabandi2018neural], PETS [chua2018deep], GPS [levine2013guided], and PILCO [deisenroth2011pilco]. Model-based approaches are more sample-efficient but may struggle in complex domains where accurate dynamics modeling is harder than direct policy learning.
@@
EAs are effective for exploring high-dimensional morphology spaces. In the outer loop, evolutionary operators determine how new morphologies are generated, thereby shaping both the breadth and depth of exploration [carvalho2024role]. In discrete design spaces, the mutation operator typically adds or removes structural elements—e.g., joints, limbs, modules, and kinematic branches—or modifies material properties. In continuous design spaces—including link lengths, joint stiffness, and mass distribution—Gaussian perturbations are commonly used. The crossover operator further expands the search space by recombining structural components or genotype encodings from multiple parents. Notably, generative models, such as Large Language Models (LLMs), have been used as high-level mutation operators to propose structured, human-interpretable edits based on prompts [song2025laser], further expanding the operator toolbox for morphology search.
@@
By alternating between individual-level controller learning and population-level morphology evolution, ERL methods reflect Darwinian evolution, enhanced by a Baldwinian learning mechanism [luo2022effects, li2023evaluation]. A representative example of this line is Neural Graph Evolution (NGE) [wang2019neural], which evolves graph-structured morphologies while training universal GNN-based controllers via PPO. NGE demonstrates that GNN controllers can generalize across various body structures, thereby reducing the cost of controller retraining during morphology search. The scalability of ERL is further illustrated by Deep Evolutionary Reinforcement Learning (DERL) [gupta2021embodied]. In contrast to NGE, DERL employs individual controllers that are trained from scratch for each morphology. Coupled with increasingly challenging environments, this framework enables the emergence of morphological intelligence, resulting in agents that exhibit increased robustness, adaptability, and behavioral diversity. Similar strategies appear in subsequent works [park2021computational, bhatia2021evolution, chen2023evolving, wang2024learn, yue2025toward, cheng2024structural, harada2024lamarckian, ringel2025text2robot, jin2025evolutionary].
@@
As summarized in Table I, recent research has expanded the classical ERL workflow along several dimensions. First, richer morphological representations have been introduced, including hyperbolic embeddings [dongleveraging2024], text-to-mesh generation [ringel2025text2robot], and latent space embedding [jeon2025convergent]. Second, advanced control-learning techniques, such as meta-RL [belmonte2022meta], transfer learning [harada2024lamarckian], imitation learning [hernandezfollowing], and policy-sharing [xie2025accelerating, li2024generating, wang2024evolution], have improved sample efficiency and generalization. A parallel line of work introduces novel morphological selection mechanisms, such as paired competition [banarse2019body], successive halving [katayama2023ga], multi-fidelity archiving [zhao2024morphological, nagiredla2024ecode], and multi-objective balancing [zhao2025multi], which can prevent premature convergence in morphology space [doncieux2014beyond, mertan2024investigating].
@@
Notably, the combination of EA and RL has also been widely explored in control optimization. These hybrid methods maintain a population of policies that evolve across generations, allowing for broad exploration of the policy space while refining individual controllers via RL’s gradient-based or value-based updates. Representative approaches include ERL [khadka2018evolution], CERL [khadka2019collaborative], PDERL [bodnar2020proximal], SERL [wang2022surrogate], and ERL-Re2 [hao2023erl]. Comprehensive reviews of this line of work can be found in [drugan2019reinforcement, sigaud2023combining, bai2023evolutionary, zhu2023survey, lin2025evolutionary, li2024bridging].
@@
Learnable (CL) Universal Hybrid (UF) Direct Predefined (Partial) Task-Driven
@@
Learnable (CL) Individual Hybrid (UF) Direct Modifiable (Full) Task-Driven
@@
LOKI [jeon2025convergent]
2025
[link]
Legged
Robot
Simulation
{MuJoCo}
Learnable (CL) Universal Hybrid (UF) Indirect Predefined (Partial) Hybrid
RL (PPO) Transformer EA (Search) Latent Locomotion Task+Diversity
@@
IV-A2 Surrogate-Assisted Methods
A central challenge in bi-level co-design arises from its hierarchical structure, where the inner-loop optimization is computationally expensive and often requires days of training for complex tasks. Surrogate-assisted methods address this issue by employing predictive models to approximate agent task performance, thereby reducing the need for repeated full-scale policy evaluations.
Bayesian Optimization (BO) [shahriari2015taking] has emerged as a prominent framework for modeling the morphology–performance relationship in ECD. As a black-box, zeroth-order optimizer, BO efficiently handles expensive objective evaluations. At each iteration, BO updates a surrogate model using previously evaluated morphologies and their performance, then selects new candidates through acquisition functions, such as Probability of Improvement (PI), Expected Improvement (EI), Entropy Search (ES), or Upper Confidence Bound (UCB), thereby balancing exploration and exploitation [wilson2018maximizing]. The selected candidates are evaluated, and their results are used to update the surrogate model for the subsequent iteration.
Gaussian Processes (GPs) are among the most commonly used surrogates in BO. For instance, HPC-BBO [liao2019data] introduced a hierarchical framework that facilitates parallel hardware validation. Similarly, [bjelonic2023learning] pre-trained a universal controller in simulation and used BO to optimize parallel–elastic spring parameters for a real quadruped, demonstrating effective sim-to-real transfer. Variants of BO with GPs have been further explored in [bhatia2021evolution, hu2021neural, saito2024effective, pan2021emergent, rajani2023co]. To enhance generalization, [chen2024pretraining] combined domain randomization with a pretraining–finetuning pipeline. Extending beyond single-task BO, [dong2024cagecoopt] proposed a multi-task BO framework for robot manipulators. More recently, [schneider2025task] combined BO with Hyperband to generate novel manipulator designs that improve performance across both seen and unseen tasks.
In addition to traditional surrogates used in BO, recent approaches have explored alternative choices. For example, Q-value functions (Q-Surro) from RL have been used to evaluate candidate morphologies [luck2020data, luck2021robot, du2025efficient]. Multi-dynamics models can further enhance generalization across different agent designs [sliacka2023co]. Alternatively, when historical interaction data is available, near-optimal control actions (A-Surro) can be partially or fully transferred from previously optimized designs to accelerate evaluation [liu2023rapidly]. More recent studies by [zhao2025cross, strgar2025accelerated] developed universal policies that span multiple tasks and can be efficiently employed as a zero-shot performance evaluator (A-Surro), thereby improving efficiency. Moreover, in [bai2025learning], manipulator morphologies are evaluated and optimized using pre-collected trajectories and a learned reward model (R-Model), with the reliance on reward model evaluations gradually increasing throughout the training process.
@@ IV-A3 Other Methods
Notably, the inner-loop controller optimization can be implemented using alternatives to RL, including trajectory optimization [fadini2022simulation], evolutionary algorithms [jelisavcic2019lamarckian, le2022morpho, le2025efficient], numerical solvers [sartore2023codesign, ha2018computational2], and information-theoretic classifiers [iii2021taskagnostic]. To accelerate bi-level co-design, Goff et al. proposed A-MEL [le2024improving], which integrates asynchronous learning and evolution to reduce both time and computational costs. Strgar et al. [Strgar-RSS-24] further leveraged massively parallel differentiable simulations to efficiently optimize neural controllers for large morphology populations.
@@ IV-B Single-Level Co-Design
As shown in Fig. 3 (b), single-level co-design treats both morphology and control as variables within a unified optimization process. We categorize recent work along this line into four subcategories: 1) EA-Based Methods, 2) RL-Based Methods, 3) Physics-Based Methods, and 4) Diversity-Based Methods, which are summarized in Table III.